OpenStudy (anonymous):

given sin(theta)=12/13, find sec(theta) a.)sec(theta)=13/12 b.)sec(theta)=13/5 c.)sec(theta)=5/13 d.)sec(theta)=5/12

OpenStudy (johnweldon1993):



OpenStudy (johnweldon1993):

So no we can see cos(theta) = 5/13 And since we know \(\large sec(\theta) = \frac{1}{cos(\theta)}\) ...

OpenStudy (anonymous):

my brain just died hun xD